what is the internal energy of the 1700 kg lead considering that its standard coefficient of thermal and specific heat that has made a 6 m cube of lead heated from 30 degrees to 300 degrees celsius


1
Expert's answer
2021-06-07T09:35:35-0400

The internal energy is given as follows:


"Q = mc\Delta T"

where "m = 1700 kg" is the mass of hte cube, "c = 1280 \space \dfrac{J}{kg\cdot \degree C}" is the specific heat capacity of lead, "\Delta T = 300\degree C - 30\degree C = 270\degree C" is the change in temperature. Thus, obtain:


"Q = 1700\cdot 1280 \cdot 270 \approx 588\times 10^6 J = 588MJ"

Answer. 588 MJ.